You have just borrowed $95,000 to buy a condo. You will repay the loan in equal monthly payments of $805 over the next 40 years. What is the effective annual rate (EAR) on the loan?

Enter your answer as a percentage rounded to two decimal places (e.g., 23.45%, NOT .2345). (What numbers do I put into the financial calculator for N, I/Y, PV, FV, and PMT?)
Answer #1

The problem can be solved on a financial calculator as described below:

- Input N = (40*12) = 480, PV = $ 95000, FV = $ 0, PMT = $ 805

- CMPT - > I/Y

- I/Y = 0.008314466 (this rate is on a per month basis)

EAR = (1.008314466)^(12) - 1 ~ 0.104465 or 10.45 %